# Using large language models for sensitivity analysis in causal inference: cases studies on Cornfield inequality and E-value

### 15 Mar 2026 v1

@@ Abstract
Sensitivity analysis methods such as the Cornfield inequality and the E-value were developed to assess the robustness of observed associations against unmeasured confounding – a major challenge in observational studies. However, the calculation and interpretation of these methods can be difficult for clinicians and interdisciplinary researchers. Recent advances in large language models (LLMs) offer accessible tools that could assist sensitivity analyses, but their reliability in this context has not been studied. We assess four widely used LLMs, ChatGPT, Claude, DeepSeek, and Gemini, on their ability to conduct sensitivity analyses using Cornfield inequalities and E-values. We first extract study-specific information (exposures, outcomes, measured confounders, and effect estimates) from four published observational studies in different fields. Using those information, we develop structured prompts to assess the performance of the LLMs in three aspects: (1) accuracy of E-value calculation, (2) qualitative interpretation of robustness to unmeasured confounding, and (3) suggestion of possible unmeasured confounders. To our knowledge, this is the first study to investigate the use of LLMs for sensitivity analysis. The results show that ChatGPT, Claude, and Gemini accurately reproduce the E-values, whereas DeepSeek shows small biases. Qualitative conclusions from all the LLMs align with the magnitude of the E-values and the reported effect sizes, and all models identify biologically and epidemiologically plausible unmeasured confounders. These findings suggest that, when guided by structured prompting, LLMs can effectively assist in evaluating unmeasured confounding, and thereby can support study design and decision-making in observational studies.
@@ 1 Introduction
For clinicians and interdisciplinary researchers who are conducting observational studies, it can be challenging to fully understand, calculate, and interpret the Cornfield inequality and the E-value. However, the advancement of Large Language Models (LLMs) could offer an accessible tool to assist with sensitivity analysis. LLMs represent a class of artificial intelligence models designed to understand and generate human-like text (Vaswani et al., 2017; Brown et al., 2020; Zhao et al., 2023). With proper input information from a study, we hypothesize that LLMs can support the calculations of the Cornfield inequality and the E-value, generate clear explanations, and suggest possible unmeasured confounders.
LLMs have increasingly been explored as tools for causal inference (Liu et al., 2025; Ma, 2025). Prior studies have primarily focused on using LLMs for tasks such as causal discovery (Jin et al., 2023b; Takayama et al., 2024; Cohrs et al., 2024; Lee et al., 2025) and causal reasoning (Yao et al., 2023; Jin et al., 2023a; Chi et al., 2024). However, using LLMs for sensitivity analysis for unmeasured confounding remains largely unexplored. To the best of our knowledge, this is the first study to develop prompting methods and to evaluate the LLMs’ ability to conduct sensitivity analyses.
In this article, we asses the performance of the LLMs in performing sensitivity analysis using case studies from four published observational studies. The studies cover topics related to smoking, back pain, Alzheimer’s disease, and environmental health. We will use the information extracted from these studies to evaluate four widely used LLMs. Our assessment is threefold: (i) quantitatively assessing if the LLMs can correctly compute the E-values, (ii) qualitatively assessing if the LLMs can generate proper conclusions on how likely the study findings are subject to unmeasured confounding, and (iii) assessing if the LLMs can suggest potential unmeasured confounders. Through this evaluation, we hope our prompting strategy may serve as a tool to help researchers better assess whether findings from observational studies are subject to unmeasured confounding, thereby helping decision-making in clinical and public health research.
@@ 2.1 Brief overview
Cornfield et al. (1959) discussed the issues against the causal role of tabacco smoking in lung cancer, and their idea was later formalized as the Cornfield inequality. They illustrated this idea with the example of smoking and lung cancer: if smokers have a ninefold higher risk of lung cancer compared to non-smokers, and if this risk is due entirely to some unmeasured factor rather than to smoking itself, then this factor would need to be extremely strongly associated with smoking (with at least a nine-fold risk ratio). However, since no such factor has been identified, the authors argued that confounding alone could not account for the observed association between smoking and lung cancer. This reasoning, which evaluates whether an unmeasured confounder could account for an observed effect, later became a core concept in sensitivity analysis for causal inference.
@@
where “RRRR” represents the risk ratio of exposure to outcome in observational studies after adjusting for measured covariates. Besides the risk ratio (RR), the E-value also applies to other effect measures, including hazard ratio (HR) and odds ratio (OR). When an obsered $RR=3.6$, the formula calculates the E-value as $3.6+\sqrt{3.6\texttimes(3.6-1)}=6.66$. Therefore, to explain away an $RR=3.6$, an unmeasured confounder would need to be associated with both exposure and outcome by at least 6.66-fold risk ratio each. A larger the E-value suggests that it is less plausible that unmeasured confounding could explain away the estimated association, and vice versa.
Figure 1 illustrates the overall workflow of our study. We select four published observational studies. From each study, we extract key information, including the exposure, outcome, measured confounders, and the estimated associations/effects. These study-specific inputs are then provided to LLMs via structured prompts. Finally, the LLMs generate the outputs of interest: calculate E-values, interpreting the robustness of the observed association to unmeasured confounding, and suggesting plausible unmeasured confounders.
2.2 Source of case study
To evaluate the performance of LLMs in sensitivity analysis, we selected four published observational studies from different fields: a smoking study, a back pain study, a Alzheimer’s study, and an environmental health study. We briefly summarize these four studies as follows:
@@ 2.2 Source of case study
  • Alzheimer’s study (Yaghmaei et al., 2024): The study investigated whether the Alzheimer’s treatment affected the five-year survival probability of Alzheimer’s disease patients.
@@
Table 1 summarizes more detailed characteristics from these four studies. All four studies clearly documented the exposures, outcomes, measured confounders, and the estimated effect sizes. In addition, they all cited VanderWeele and Ding (2017), and calculated and reported the E-values. Since some studies included multiple exposures, a total of 11 E-values were calculated across the four studies. Except for the Alzheimer’s study, other three studies explicitly made conclusions and interpretations regarding the level of unmeasured confounding based on the calculated E-values.
Table 1: Summary characteristics of the four observatianal studies. The exposures, outcomes, measured confounders, and effect sizes in those paper will used as the input to the large language models. BMI: body mass index.
2.3 Analysis to assess the performance of LLM on sensitivity analysis
@@
We evaluate four popular LLMs: ChatGPT-5 mini (Achiam et al., 2023), Claude 4.5 Opus (Anthropic, 2025), Deepseek V3 (Liu et al., 2024), and Gemini-2.5-pro (Team et al., 2023). We develop a thorough prompt strategy to guide those LLMs for sensitivity analysis. The full prompts are included in Supplementary File Section S1. Our prompt method begins by using an initialization prompt to provide the LLMs with relevant context:
@@ 1. Calculate the E-value using the appropriate formula
Finally, as an exploratory evaluation, we let the LLMs suggest for potential unmeasured confounders relevant to the input study. This task was intended to assess whether LLMs could recognize possible sources of confounding beyond the variables reported in the original studies. Such suggestions can help researchers better identify important variables to measure in the design of their studies.
@@ 2.4 Implementation
As proprietary LLMs cannot be deployed locally, all inference is conducted through the APIs of the LLMs. For the model’s hyperparameter configuration, to optimize accuracy and reproducibility, we set the temperature to 0 to reduce output randomness and suppress hallucination generation. Meanwhile, to accommodate long text outputs containing all reasoning steps, we set the maximum number of generated tokens to 2000. All inputs and outputs were processed through Python scripts to ensure consistency in format, token budget, and output structure across models. The codes are included in a GitHub repository https://github.com/qingyan16/LLMs-for-sensitivity-analysis.
@@ 3.1 LLMs’ quantitative ability in calculating E-values
This subsection evaluates the quantitative ability of LLMs in calculating the E-values (Table 2) based on the summarized study information. Since some studies included multiple exposures-outcome pairs, a total of 11 E-value calculations are generated. The calculated E-values from ChatGPT, Claude, and Gemini exactly match the ground truth E-values reported in the original studies, showing zero bias across all cases. However, DeepSeek shows biases ranging from -0.01 to 0.23 from the true E-values. Notably, although the explicit formula for calculating the E-value is not provided in the prompts, ChatGPT, Claude, and Gemin can still accurately calculate E-values using summarized information of exposures, outcomes, measured confounders, and effect estimates.
@@
Table 3 shows the qualitative conclusions generated by the LLMs based on both the Cornfield inequality and the E-value. Detailed reasoning outputs from both perspectives are provided in Supplementary File Section S2. From the reasoning outputs, the LLMs demonstrate the ability to apply the Cornfield inequality properly. For example, in the smoking study, the LLMs translate every observed effect size into the required strength of association that a unmeasured confounder would need to have with both the exposure and the outcome. In particular, the reasoning from ChatGPT and Claude varies with the magnitude of the effect size. For a larger effect (RR = 2.132) in the smoking study, both models respond that it is “less plausible” of a single unmeasured confounder explaining away the association, while for smaller effects (e.g., RR = 1.259), both models conclude such unmeasured confounding is plausible.
@@
3.3 Potential unmeasured confounders suggested by the LLMs. Each LLM are required to suggest top 3 unmeasured confounders.
@@ 4 Discussion
Our work investigates the performance of four widely used large language models (ChatGPT, Claude, DeepSeek, and Gemini) in conducting sensitivity analyses. Using studies from four different clinical/epidemiological domains, we provided the LLMs with summarized study information including the exposures, outcomes, measured confounders, and effect estimates. We developed structured prompts to guide the LLMs to perform different tasks for sensitivity analysis. Overall, the results show that ChatGPT, Claude, and Gemini can accurately calculate E-values, properly assess the robustness to unmeasured confounding, and identify reasonable potential unmeasured confounders.
The conclusions generated by the LLMs regarding how likely unmeasured confounding could explain away the observed associations are appropriately differentiated. In contrast, the original observational studies typically provide a single qualitative interpretation, even when multiple exposure-outcome cases are reported. For example, in the smoking study, although the reported effect sizes range from 1.259 to 2.132, the authors drew a single conclusion that it is not plausible for unmeasured confounding to explain away the observed effects. However, the ChatGPT conclusions distinguish among exposure-outcome pairs, suggesting that unmeasured confounding is “unlikely” when the effect size is 2.132 but “possibly” when the effect size is 1.259. These conclusions, which are based on both the Cornfield inequality and the E-value, suggest that LLMs can generate interpretations that align with the magnitude of effect estimates and corresponding E-values.
As this is a case study, in this work, we include only four published studies as input for the LLMs. In future research, we could use a larger number of observational studies that apply the Cornfield inequality and the E-value, allowing for a more thorough evaluation.
@@
Acknowledgement
Authors would like to thank the journal Observational Studies for their announcement of the special issue on Cornfield inequality and sensitivity analyses, which motivated this research.

### 09 Apr 2026 v3

# Using large language models for sensitivity analysis in causal inference: case studies on Cornfield inequality and E-value

@@ Abstract
Sensitivity analysis methods such as the Cornfield inequality and the E-value were developed to assess the robustness of observed associations against unmeasured confounding – a major challenge in observational studies. However, the calculation and interpretation of these methods can be difficult for clinicians and interdisciplinary researchers. Recent advances in large language models (LLMs) offer accessible tools that could assist sensitivity analyses, but their reliability in this context has not been studied. We assess four widely used LLMs, ChatGPT, Claude, DeepSeek, and Gemini, on their ability to conduct sensitivity analyses using Cornfield inequalities and E-values. We first extract study-specific information (exposures, outcomes, measured confounders, and effect estimates) from four published observational studies in different fields. Using such information, we develop structured prompts to assess the performance of the LLMs in three aspects: (1) accuracy of E-value calculation, (2) qualitative interpretation of robustness to unmeasured confounding, and (3) suggestion of possible unmeasured confounders. To our knowledge, there has been little prior work on using LLMs for sensitivity analysis, and this study is an early investigation in this area. The results show that ChatGPT, Claude, and Gemini accurately reproduce the E-values, whereas DeepSeek shows small biases. Qualitative conclusions from all the LLMs align with the magnitude of the E-values and the reported effect sizes, and all models identify biologically and epidemiologically plausible unmeasured confounders. These findings suggest that, when guided by structured prompts, LLMs can effectively assist in evaluating unmeasured confounding, and thereby can support study design and decision-making in observational studies.
@@ 1 Introduction
For clinicians and interdisciplinary researchers who are conducting observational studies, it can be challenging to fully understand, calculate, and interpret the Cornfield inequality and the E-value. However, advances in large language models (LLMs) offer accessible tools that could assist with sensitivity analyses. LLMs represent a class of artificial intelligence models designed to understand and generate human-like text (Vaswani et al., 2017; Brown et al., 2020; Zhao et al., 2023). With proper input information from a study, we hypothesize that LLMs can support the calculations of the Cornfield inequality and the E-value, generate clear explanations, and suggest possible unmeasured confounders.
LLMs have increasingly been explored as tools for causal inference (Liu et al., 2025; Ma, 2025). Prior studies have primarily focused on using LLMs for tasks such as causal discovery (Jin et al., 2023b; Takayama et al., 2024; Cohrs et al., 2024; Lee et al., 2025) and causal reasoning (Yao et al., 2023; Jin et al., 2023a; Chi et al., 2024). However, using LLMs for sensitivity analysis for unmeasured confounding remains largely unexplored, and to the best of our knowledge, this study is an early investigation of this topic.
In this article, we assess the performance of the LLMs in performing sensitivity analysis using case studies from four published observational studies. The studies cover topics related to smoking, back pain, Alzheimer’s disease, and environmental health. We will use the information extracted from these studies to evaluate four widely used LLMs. Our assessment is threefold: (i) quantitatively assessing if the LLMs can correctly compute the E-values, (ii) qualitatively assessing if the LLMs can generate proper conclusions on how likely the study findings are subject to unmeasured confounding, and (iii) assessing if the LLMs can suggest potential unmeasured confounders. Through this evaluation, we hope our prompting strategy can serve as a tool to help researchers better assess whether findings from observational studies are subject to unmeasured confounding, thereby helping decision-making in clinical and public health research.
@@ 2.1 Brief overview
Cornfield et al. (1959) discussed the issues against the causal role of tobacco smoking in lung cancer, and their idea was later formalized as the Cornfield inequality. They illustrated this idea with the example of smoking and lung cancer: if smokers have a ninefold higher risk of lung cancer compared to non-smokers, and if this risk is due entirely to some unmeasured factor rather than to smoking itself, then this factor would need to be extremely strongly associated with smoking (with at least a nine-fold risk ratio). However, since no such factor has been identified, the authors argued that confounding alone could not account for the observed association between smoking and lung cancer. This reasoning, which evaluates whether an unmeasured confounder could account for an observed effect, later became a core concept in sensitivity analysis for causal inference.
@@
where “RRRR” represents the risk ratio of exposure to outcome in observational studies after adjusting for measured covariates. Besides the risk ratio ($R$), the E-value also applies to other effect measures, including hazard ratio (HR) and odds ratio (OR). When an observed $RR=3.6$, the formula calculates the E-value as $3.6+\sqrt{3.6\texttimes(3.6-1)}=6.66$. Therefore, to explain away an $RR=3.6$, an unmeasured confounder would need to be associated with both exposure and outcome by at least 6.66-fold risk ratio each. A larger E-value suggests that it is less plausible that unmeasured confounding could explain away the estimated association, and vice versa.
Figure 1 illustrates the overall workflow of our study. We select four published observational studies. From each study, we extract key information, including the exposure, outcome, measured confounders, and the estimated associations/effects. These study-specific inputs are then provided to LLMs via structured prompts. Finally, the LLMs generate the outputs of interest: calculating E-values, interpreting the robustness of the observed association to unmeasured confounding, and suggesting plausible unmeasured confounders.
2.2 Sources of case studies
To evaluate the performance of LLMs in sensitivity analysis, we select four published observational studies from different fields: a smoking study, a back pain study, an Alzheimer’s study, and an environmental health study. We briefly summarize these four studies as follows:
@@ 2.2 Sources of case studies
  • Alzheimer’s study (Yaghmaei et al., 2024): The study investigated whether Alzheimer’s treatment affected the five-year survival probability of Alzheimer’s disease patients.
@@
Table 1 summarizes more detailed characteristics from these four studies. All four studies clearly documented the exposures, outcomes, measured confounders, and estimated effect sizes. In addition, they all cited VanderWeele and Ding (2017), and calculated and reported the E-values. Since some studies included multiple exposures, a total of 11 E-values were calculated across the four studies. Except for the Alzheimer’s study, the other three studies explicitly made conclusions and interpretations regarding the level of unmeasured confounding based on the calculated E-values.
Table 1: Summary characteristics of the four observational studies. The exposures, outcomes, measured confounders, and effect sizes in those paper will be used as the input to the large language models. BMI: body mass index.
2.3 Analysis to assess the performance of LLM in sensitivity analysis
@@
We evaluate four popular LLMs: ChatGPT-5 mini (Achiam et al., 2023), Claude 4.5 Opus (Anthropic, 2025), DeepSeek V3 (Liu et al., 2024), and Gemini-2.5-pro (Team et al., 2023). We develop a thorough prompt strategy to guide those LLMs for sensitivity analysis. The full prompts are included in Section S1 of the supplementary file, which is available in the GitHub repository of the authors. Our prompt method begins by using an initialization prompt to provide the LLMs with relevant context:
@@ 1. Calculate the E-value using the appropriate formula
Finally, as an exploratory evaluation, we let the LLMs suggest potential unmeasured confounders relevant to the input study. This task is intended to assess whether LLMs could recognize possible sources of confounding beyond the variables reported in the original studies. Such suggestions can help researchers better identify important variables to measure in the design of their studies.
@@ 2.4 Implementation
As proprietary LLMs cannot be deployed locally, all inference is conducted through the APIs of the LLMs. To optimize accuracy and reproducibility, we set the temperature to 0 to reduce output randomness and suppress hallucinated content. Meanwhile, to accommodate long text outputs containing all reasoning steps, we set the maximum number of generated tokens to 2000. All inputs and outputs are processed through Python scripts to ensure consistency in format, token budget, and output structure across models. The codes are included in a GitHub repository https://github.com/qingyan16/LLMs-for-sensitivity-analysis.
@@ 3.1 LLMs’ quantitative ability in calculating E-values
This subsection evaluates the quantitative ability of LLMs in calculating the E-values (Table 2) based on the summarized study information. Since some studies included multiple exposure-outcome pairs, a total of 11 E-value calculations are generated. The calculated E-values from ChatGPT, Claude, and Gemini exactly match the ground truth E-values reported in the original studies, showing zero bias across all cases. However, DeepSeek shows biases ranging from -0.01 to 0.23 relative to the true E-values. Notably, although the explicit formula for calculating the E-value is not provided in the prompts, ChatGPT, Claude, and Gemin can still accurately calculate E-values using summarized information of exposures, outcomes, measured confounders, and effect estimates.
@@
Table 3 shows the qualitative conclusions generated by the LLMs based on both the Cornfield inequality and the E-value. Detailed reasoning outputs from both perspectives are provided in Supplementary File Section S2. From the reasoning outputs, the LLMs demonstrate the ability to apply the Cornfield inequality properly. For example, in the smoking study, the LLMs translate every observed effect size into the required strength of association that an unmeasured confounder would need to have with both the exposure and the outcome. In particular, the reasoning from ChatGPT and Claude varies with the magnitude of the effect size. For a larger effect ($R$ = 2.132) in the smoking study, both models respond that it is “less plausible” of a single unmeasured confounder explaining away the association, while for smaller effects (e.g., $R$ = 1.259), both models conclude such unmeasured confounding is plausible.
@@
3.3 Potential unmeasured confounders suggested by the LLMs. Each LLM is required to suggest top 3 unmeasured confounders.
@@ 4 Discussion
Our work investigates the performance of four widely used large language models (ChatGPT, Claude, DeepSeek, and Gemini) in conducting sensitivity analyses. Using studies from four different clinical and epidemiological domains, we provide the LLMs with summarized study information including the exposures, outcomes, measured confounders, and effect estimates. We develop structured prompts to guide the LLMs to perform different tasks for sensitivity analysis. Overall, the results show that ChatGPT, Claude, and Gemini can accurately calculate E-values, properly assess robustness to unmeasured confounding, and identify reasonable potential unmeasured confounders.
The conclusions generated by the LLMs regarding how likely unmeasured confounding could explain away the observed associations are appropriately differentiated. In contrast, the original observational studies typically provide a single qualitative interpretation, even when multiple exposure-outcome cases are reported. For example, in the smoking study, although the reported effect sizes range from 1.259 to 2.132, this article drew a single conclusion that it is not plausible for unmeasured confounding to explain away the observed effects. However, the ChatGPT conclusions distinguish among exposure-outcome pairs, suggesting that unmeasured confounding is “unlikely” when the effect size is 2.132 but “possibly” when the effect size is 1.259. These conclusions, which are based on both the Cornfield inequality and the E-value, suggest that LLMs can generate interpretations that align with the magnitude of effect estimates and corresponding E-values.
As this is a case study, we include only four published studies as inputs to the LLMs. In future research, we could use a larger number of observational studies that apply the Cornfield inequality and the E-value, allowing for a more thorough evaluation.
@@
Acknowledgements
The authors would like to thank the journal Observational Studies for their announcement of the special issue on Cornfield inequality and sensitivity analyses, which motivated this research.